# Friend or Foe Inside? Exploring In-Process Isolation to Maintain Memory Safety for Unsafe Rust

### v1

@@ II-B Process-based Isolation
Under process-based isolation, for different compartments (i.e., processes) to communicate and exchange data with each other, they must do so via inter-process communication (IPC) primitives. IPC typically comes with high overheads due to the context switching associated with scheduling different process and data crossing security and failure boundaries. In Sandcrust, argument passing and communicating return values between the main process and sandboxed code is handled over IPC channels. To allow the passing of Rust (and C) objects across the process boundary, one must be able to serialize, and deserialize, any data type that is to be transferred. As we will show in Section IV-D the overhead associated with serialization and deserialization is the principle source of run-time overhead for compartmentalization in Sandcrust. Consequently, minimizing the overhead of object passing to and from the sandboxed code is an important consideration for making automatic program compartmentalization in Rust more tractable.
@@ II-C In-Process Isolation
Automatic compartmentalization of multi-language Rust applications using PKU has been explored by Liu et al. [6] and Kirth et al. [7]. These prior works focus on portioning the heap into distinct protection domains for safe and unsafe Rust code, including calls occurring via FFIs. While these compartmentalization approaches share similar goals to process-based isolation, they are not directly comparable as process-based isolation encompasses not only the application heap, but all data, including the stack and static allocations. As discussed in Section II-B, process-based isolation can further provide a degree of resilience against memory corruption whereas conventional SFI approaches terminate the application as soon as a violation of a domain boundary is detected.
@@ III Problem Statement
The validation phase then compares the resulting prototype with Sandcrust as a representative of automated process isolation for the Rust FFI. In particular, we look at the performance overhead of the two solutions, micro-benchmarking the context switch cost as well as evaluating real-world applications for different data transfer volumes. Moreover we compare the security posture of the two solutions. All study results are discussed in Section VI along with lessons learned.
@@ IV-B SDRaD-FFI design
The sandbox! macro is expanded into calls to the SDRaD library and additional glue logic for passing parameters and results between the domains as well as error handling. Syntactically, the definition of a sandboxed function that takes a number of arguments and returns some value is maintained as __real_F(), while an additional wrapper __wrap_F() is defined to execute the required glue logic in the nested domain. Figure 1 shows the general control flow for executing a call to F in a nested domain .
@@
the function arguments (cf. Section IV-D) and writes them to a
@@
(cf. Section IV-D). Then, __real_F() holding the original function
@@ IV-C In-Process Communication
Some of the operations above are not natively supported by the SDRaD API, which we extended (cf. Section IV-E).
@@ IV-E SDRaD Integration and Extension
As we discussed in Section IV-C, transferring data between domains requires that the nested domain stack information should be exposed to the parent domain. We extend the SDRaD API with two new APIs: 1) exposing the stack base pointer to the caller, allowing the parent domain to write argument vector metadata in the nested domain stack area. 2) updating the stack base pointer of the nested domain that prevents overwriting information stored on the nested domain’s stack by the parent domain. These extensions were sufficient to support the in-process communication mechanism described above.
@@ V-A Microbenchmark
To gauge the overall context switch costs, we sandboxed the empty() function and measured its execution time, i.e., the time for entering and exiting the sandbox. We compared SDRaD-FFI to a baseline value without compartmentalization, Sandcrust, and PKRU-Safe. Note that the PKRU-Safe microbenchmark value is an estimate based on the description in the paper [7]. Moreover, PKRU-Safe only compartmentalizes heap memory space, and we report the number here as a reference for another PKU-based isolation mechanism. Figure 2 shows the execution times for all systems under test. While the sandboxed empty() function with SDRaD-FFI takes (=), the baseline without sandboxing takes (=) and Sandcrust takes (=). SDRaD-FFI is much faster than Sandcrust and slower than the baseline function calls. This result is in the same order of magnitude as context switch overheads reported for PKRU-Safe [7].
@@ V-B Snappy
To optimize Abomonation for , we used the unstable specialization feature as explained Section IV-D. Our experiments show that this specialization improved performance significantly as seen in Figure 2(a).
@@
Overall, as described in Section IV-D, the memory layout presentation used by Abomonation to serialize and deserialize variables is suitable for optimization, unlike Bincode, which uses binary format representation and cannot be optimized automatically by the compiler in the same way.
@@
We perform the following benchmarking using the original version of Abomonation. Figure 3 show the snappy compress() and uncompress() function execution times for different byte sizes: 1 byte to bytes. In Figure 2(b), Sandcrust performs worst when compressing data smaller than bytes which can be explained by our microbenchmark results from Section V-A.
@@ V-C Libpng
Table I shows mean and relative-standard deviation for decoding measurement of different images. SDRaD-FFI introduced a worst-case overhead of performance degradation compared to a baseline measurement for image decoding and it performed significantly better than Sandcrust in all cases.
@@ VI-A Lessons Learned
Prior work on automated software compartmentalization conventionally relies on compiler-based source-to-source translation which, as discussed in Section II-B, faces a high barrier for adoption due to requiring invasive changes to compilers and toolchains. Compartmentalization APIs such as SDRaD place the burden of integrating compartmentalization capability on the developers. For languages such as C, such changes can become invasive as the amount of data to be passed across the domain boundaries increases. Languages with strong metaprogramming capabilities, such as the Rust macro system can greatly reduce the complexity for repetitive integration effort, such as argument passing. There can also be benefits in converging on interoperable interfaces for language- and library-level compartmentalizing; by adopting a similar interface as Sandcrust, SDRaD-FFI can be used as a drop-in replacement for Sandcrust that provides improved run-time performance.
@@ VII Related Work
Several studies have explored the compartmentalizing of applications based on different underlying primitives: hardware-assisted in-process isolation [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 7, 28, 29, 9], process-based isolation [5] and software-based isolation [12, 6]. The majority of such approaches address the problem of compartmentalizing applications written in C or C++ [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 9]. In this work, we focus on multi-language applications, specifically Rust code calling C or C++ code [11, 8, 7, 5, 6]. As noted in Section II, the majority of prior work on using in-process isolation operate under weaker isolation guarantees than those of comparable solutions based on process-isolation [5]; XRust [6], Galeed [54] and PKRU-Safe [7] only protect the Rust heap from unsafe code while Fidelius Charm [11] only protects the Rust stack.
@@
Independently and concurrently with our work, Bang et al. [8] propose TRust, a scheme with similar goals as Sandcrust [5]. TRust protects both the Rust heap and stack from unsafe and mixed-language code. However, unlike SDRaD-FFI, which we designed to act as a drop-in replacement for Sandcrust, TRust relies on compiler-driven compartmentalization of Rust code based on the language-level boundary between safe and unsafe Rust. The principal benefit of compiler-based approaches, such as PKRU-Safe and TRust is relieving the developer completely of the burden of defining protection domains. In practice, automated approaches are restricted to boundaries that can be derived from language-level constructs. In the case of Rust, approaches such as TRust are not suitable for isolating unsafe code from other unsafe code. Combined with conservative compartmentalization policies that classify any objects that the untrusted code might use as unsafe, this can result in problematic edge cases where nearly all allocations are delegated to the unsafe domain (cf. Table 10 in [8]), negating the benefit of compartmentalization in the first place. Approaches, such as Sandcrust and SDRaD-FFI, which minimize developer effort but leave the developer in control can deal efficiently with such cases. Fully automating compartmentalization using compiler-based rewriting can also, at least in the short term, be counter-productive for the adoption of these solutions as invasive changes to mature toolchains are required, which are unlikely to make their way upstream. Point-solutions that can be deployed as discrete libraries or crates offer a more realistic path to industry adoption. Finally, as discussed in Section II, one of the principal motivations for process-based isolation are improved resilience by means of a failure boundary between processes. Apart from SDRaD and SDRaD-FFI, no prior work in in-process isolation provides such a failure boundary between compartmentalized domains.
@@
Table II summarizes our comparison between the relevant compartmentalization schemes for commodity 64-bit x86 processors excluding approaches that require kernel modification [11, 20, 21] for deployment. SDRaD-FFI is the only in-process isolation solution that provides similar capabilities as process-based isolation while significantly improving the run-time overhead compared to the similar Sandcrust in the snappy use case. While SDRaD-FFI remains less efficient than in-process isolation with weaker security guarantees such as TRust, we point out that the measured performance overhead is highly benchmark-specific. We posit that schemes such as SDRaD-FFI remain a valid alternative for compiler-based rewriting for use cases that require more flexibility than static analysis can achieve to take developer intent into account.

### v2

@@ II-B Process-based Isolation
Under process-based isolation, for different compartments (i.e., processes) to communicate and exchange data with each other, they must do so via inter-process communication (IPC) primitives. IPC typically comes with high overheads due to the context switching associated with scheduling different process and data crossing security and failure boundaries. In Sandcrust, argument passing and communicating return values between the main process and sandboxed code is handled over IPC channels. To allow the passing of Rust (and C) objects across the process boundary, one must be able to serialize, and deserialize, any data type that is to be transferred. As we will show in Section˜IV-D the overhead associated with serialization and deserialization is the principle source of run-time overhead for compartmentalization in Sandcrust. Consequently, minimizing the overhead of object passing to and from the sandboxed code is an important consideration for making automatic program compartmentalization in Rust more tractable.
@@ II-C In-Process Isolation
Automatic compartmentalization of multi-language Rust applications using PKU has been explored by Liu et al. [6] and Kirth et al. [7]. These prior works focus on portioning the heap into distinct protection domains for safe and unsafe Rust code, including calls occurring via FFIs. While these compartmentalization approaches share similar goals to process-based isolation, they are not directly comparable as process-based isolation encompasses not only the application heap, but all data, including the stack and static allocations. As discussed in Section˜II-B, process-based isolation can further provide a degree of resilience against memory corruption whereas conventional SFI approaches terminate the application as soon as a violation of a domain boundary is detected.
@@ III Problem Statement
The validation phase then compares the resulting prototype with Sandcrust as a representative of automated process isolation for the Rust FFI. In particular, we look at the performance overhead of the two solutions, micro-benchmarking the context switch cost as well as evaluating real-world applications for different data transfer volumes. Moreover we compare the security posture of the two solutions. All study results are discussed in Section˜VI along with lessons learned.
@@ IV-B SDRaD-FFI design
The sandbox! macro is expanded into calls to the SDRaD library and additional glue logic for passing parameters and results between the domains as well as error handling. Syntactically, the definition of a sandboxed function that takes a number of arguments and returns some value is maintained as __real_F(), while an additional wrapper __wrap_F() is defined to execute the required glue logic in the nested domain. Figure˜1 shows the general control flow for executing a call to F in a nested domain .
@@
the function arguments (cf. Section˜IV-D) and writes them to a
@@
(cf. Section˜IV-D). Then, __real_F() holding the original function
@@ IV-C In-Process Communication
Some of the operations above are not natively supported by the SDRaD API, which we extended (cf. Section˜IV-E).
@@ IV-E SDRaD Integration and Extension
As we discussed in Section˜IV-C, transferring data between domains requires that the nested domain stack information should be exposed to the parent domain. We extend the SDRaD API with two new APIs: 1) exposing the stack base pointer to the caller, allowing the parent domain to write argument vector metadata in the nested domain stack area. 2) updating the stack base pointer of the nested domain that prevents overwriting information stored on the nested domain’s stack by the parent domain. These extensions were sufficient to support the in-process communication mechanism described above.
@@ V-A Microbenchmark
To gauge the overall context switch costs, we sandboxed the empty() function and measured its execution time, i.e., the time for entering and exiting the sandbox. We compared SDRaD-FFI to a baseline value without compartmentalization, Sandcrust, and PKRU-Safe. Note that the PKRU-Safe microbenchmark value is an estimate based on the description in the paper [7]. Moreover, PKRU-Safe only compartmentalizes heap memory space, and we report the number here as a reference for another PKU-based isolation mechanism. Figure˜2 shows the execution times for all systems under test. While the sandboxed empty() function with SDRaD-FFI takes (=), the baseline without sandboxing takes (=) and Sandcrust takes (=). SDRaD-FFI is much faster than Sandcrust and slower than the baseline function calls. This result is in the same order of magnitude as context switch overheads reported for PKRU-Safe [7].
@@ V-B Snappy
To optimize Abomonation for , we used the unstable specialization feature as explained Section˜IV-D. Our experiments show that this specialization improved performance significantly as seen in Figure˜3(a).
@@
Overall, as described in Section˜IV-D, the memory layout presentation used by Abomonation to serialize and deserialize variables is suitable for optimization, unlike Bincode, which uses binary format representation and cannot be optimized automatically by the compiler in the same way.
@@
We perform the following benchmarking using the original version of Abomonation. Figure˜3 show the snappy compress() and uncompress() function execution times for different byte sizes: 1 byte to bytes. In Figure˜3(b), Sandcrust performs worst when compressing data smaller than bytes which can be explained by our microbenchmark results from Section˜V-A.
@@ V-C Libpng
Table˜I shows mean and relative-standard deviation for decoding measurement of different images. SDRaD-FFI introduced a worst-case overhead of performance degradation compared to a baseline measurement for image decoding and it performed significantly better than Sandcrust in all cases.
@@ VI-A Lessons Learned
Prior work on automated software compartmentalization conventionally relies on compiler-based source-to-source translation which, as discussed in Section˜II-B, faces a high barrier for adoption due to requiring invasive changes to compilers and toolchains. Compartmentalization APIs such as SDRaD place the burden of integrating compartmentalization capability on the developers. For languages such as C, such changes can become invasive as the amount of data to be passed across the domain boundaries increases. Languages with strong metaprogramming capabilities, such as the Rust macro system can greatly reduce the complexity for repetitive integration effort, such as argument passing. There can also be benefits in converging on interoperable interfaces for language- and library-level compartmentalizing; by adopting a similar interface as Sandcrust, SDRaD-FFI can be used as a drop-in replacement for Sandcrust that provides improved run-time performance.
@@ VII Related Work
Several studies have explored the compartmentalizing of applications based on different underlying primitives: hardware-assisted in-process isolation [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 7, 28, 29, 9], process-based isolation [5] and software-based isolation [12, 6]. The majority of such approaches address the problem of compartmentalizing applications written in C or C++ [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 9]. In this work, we focus on multi-language applications, specifically Rust code calling C or C++ code [11, 8, 7, 5, 6]. As noted in Section˜II, the majority of prior work on using in-process isolation operate under weaker isolation guarantees than those of comparable solutions based on process-isolation [5]; XRust [6], Galeed [54] and PKRU-Safe [7] only protect the Rust heap from unsafe code while Fidelius Charm [11] only protects the Rust stack.
@@
Independently and concurrently with our work, Bang et al. [8] propose TRust, a scheme with similar goals as Sandcrust [5]. TRust protects both the Rust heap and stack from unsafe and mixed-language code. However, unlike SDRaD-FFI, which we designed to act as a drop-in replacement for Sandcrust, TRust relies on compiler-driven compartmentalization of Rust code based on the language-level boundary between safe and unsafe Rust. The principal benefit of compiler-based approaches, such as PKRU-Safe and TRust is relieving the developer completely of the burden of defining protection domains. In practice, automated approaches are restricted to boundaries that can be derived from language-level constructs. In the case of Rust, approaches such as TRust are not suitable for isolating unsafe code from other unsafe code. Combined with conservative compartmentalization policies that classify any objects that the untrusted code might use as unsafe, this can result in problematic edge cases where nearly all allocations are delegated to the unsafe domain (cf. Table 10 in [8]), negating the benefit of compartmentalization in the first place. Approaches, such as Sandcrust and SDRaD-FFI, which minimize developer effort but leave the developer in control can deal efficiently with such cases. Fully automating compartmentalization using compiler-based rewriting can also, at least in the short term, be counter-productive for the adoption of these solutions as invasive changes to mature toolchains are required, which are unlikely to make their way upstream. Point-solutions that can be deployed as discrete libraries or crates offer a more realistic path to industry adoption. Finally, as discussed in Section˜II, one of the principal motivations for process-based isolation are improved resilience by means of a failure boundary between processes. Apart from SDRaD and SDRaD-FFI, no prior work in in-process isolation provides such a failure boundary between compartmentalized domains.
@@
Table˜II summarizes our comparison between the relevant compartmentalization schemes for commodity 64-bit x86 processors excluding approaches that require kernel modification [11, 20, 21] for deployment. SDRaD-FFI is the only in-process isolation solution that provides similar capabilities as process-based isolation while significantly improving the run-time overhead compared to the similar Sandcrust in the snappy use case. While SDRaD-FFI remains less efficient than in-process isolation with weaker security guarantees such as TRust, we point out that the measured performance overhead is highly benchmark-specific. We posit that schemes such as SDRaD-FFI remain a valid alternative for compiler-based rewriting for use cases that require more flexibility than static analysis can achieve to take developer intent into account.